OpenStudy (anonymous):

Simplify by factoring. Assume all expressions under radicals represent nonnegative numbers. √x^14 Exact answer, using radicals as needed.

OpenStudy (anonymous):

it wouldn't just be x^7?

jimthompson5910 (jim_thompson5910):

\[\large \sqrt{x^{14}}=x^{\frac{14}{2}}=x^7\]